# HGRN2: Gated Linear RNNs with State Expansion

### v1

@@ 4.1 MQAR
Multi-Query Associative Recall (MQAR) (Arora et al., 2023) is an enhanced version of the synthetic induction head dataset (Fu et al., 2023) to test the in-context associative recall ability for subquadractic models. Arora et al. (2023) found strong correlations between the accuracy of MQAR and the performance of language modeling. Our experimental setting strictly follows the original paper  [^5] https://github.com/HazyResearch/zoology, other technical details can also be found in the paper. Our hyperparameter sweeping range is: expand ratio $\in\{64,128\}$, learning rate $\in\{1e-5,5e-5,1e-4,5e-4,1e-3,5e-3,1e-2\}$.
@@
We can see from Fig.3 that HGRN2 significantly outperforms HGRN1 across various model dimensions, showing the benefit of using state expansion.
Figure 3: Results on MQAR, where the x-axis represents the model dimension and the y-axis represents accuracy, with the task becoming more difficult as the sequence length increases. HGRN2 outperforms HGRN1 in all scenarios.
@@ 4.2 Language modeling
Table 5: Results on Wikitext-103. $\downarrow$ means lower is better. We adopted the configuration of HGRN1 44m for wikitext, and it can be seen that HGRN2 significantly outperforms HGRN1 as well as all other methods.
Model PPL (val)$\downarrow$ PPL (test)$\downarrow$ Params (M)
Attn-based
@@ 4.2.1 Wiki103
For the Wikitext-103 experiment, we followed the configuration of HGRN1 to validate the performance of 44M models against a wide range of subquadractic models: FLASH (Hua et al., 2022), 1+elu (Katharopoulos et al., 2020), Performer (Choromanski et al., 2021), cosFormer (Qin et al., 2022b), Syn(D), Syn(R) (Tay et al., 2021a), gMLP (Liu et al., 2021), S4 (Gu et al., 2022a), DSS (Gupta et al., 2022b), RWKV-v4 (Peng et al., 2023), LRU (Orvieto et al., 2023), HGRN1 (Qin et al., 2023c), TNN (Qin et al., 2023a).
@@
Table 5 shows the result. We can see that HGRN2 achieves the lowest perplexity among all compared subquadractic models. In particular, HGRN2 clearly outperforms HGRN1 using even fewer parameters.
Figure 4: We compared the loss curves of HGRN2, LLaMA, Mamba, and HGRN1 under the consumption of 30 billion tokens. HGRN2 is comparable to the state-of-the-art (SOTA) methods at 1 billion and better than SOTA methods at 3 billion size.
@@ 4.2.2 The Pile
Setting.
We conduct a controlled experiment in a relatively large-scale setting on the Pile (Gao et al., 2020) to compare the LLaMa architecture (i.e., Transformer++), Mamba (Gu & Dao, 2023), HGRN1 and HGRN2. Due to the limiation of computational resources, we only train the 1B/3B models for 30B tokens.
Result.
The training curve is shown in Fig. 4. We can see that, for 1B models, HGRN2 slightly underperforms Mamba. For 3B models, HGRN2 outperforms both LLaMa and Mamba, indicating the potential to further scale up HGRN2.
4.2.3 Downstream evaluation
We train 150M/350M/1B/3B HGRN2 for 100B tokens sampled from subsets within the Pile, C4 and Wikipedia datasets and evaluate them in Commonsense reasoning tasks. Table 6 shows the result. HGRN2 outperforms HGRN1 in almost all benchmarks and competes strongly against other models.
Table 6: Performance Comparison on Commonsense Reasoning Tasks. PS, T, HS, WG stand for parameter size (billion), tokens (billion), HellaSwag, and WinoGrande, respectively. $\dagger$ indicates the results reported by Yang et al. (2023)
Model Params Token PIQA HS WG ARC-e ARC-c OBQA AVG
GPT-Neo 0.13 300 63.06 30.40 50.43 43.73 23.12 26.20 39.49
OPT 0.16 300 62.95 31.35 50.43 43.52 22.70 28.00 39.83
Pythia 0.16 300 61.32 30.16 51.93 43.18 23.12 26.80 39.42
RWKV-4 0.17 - 65.07 32.26 50.83 47.47 24.15 29.60 41.56
@@ 4.2.3 Downstream evaluation
BLOOM 0.56 350 64.09 36.97 52.80 47.35 23.98 28.20 42.23
RWKV-4 0.43 - 67.52 40.90 51.14 52.86 25.17 32.40 45.00
HGRN1 0.35 100 66.70 38.12 51.70 49.20 25.26 30.60 43.60
HGRN2 0.39 100 69.97 46.16 52.72 53.58 23.98 32.40 46.47
@@
${}^{\dagger}$ 1.3 100 71.90 49.90 54.40 61.80 28.40 34.20 50.10
${}^{\dagger}$ 1.3 100 71.40 51.20 54.10 60.10 30.40 33.80 50.20
${}^{\dagger}$ 1.3 100 69.80 47.50 51.10 58.50 27.40 31.80 47.68
HGRN1 1.0 100 70.89 48.02 51.62 55.64 27.90 31.60 47.61
HGRN2 1.0 100 74.16 54.85 56.12 58.71 27.22 34.00 50.84
OPT 2.7 300 73.83 60.60 61.01 60.77 31.31 35.2.0 53.79
@@
BLOOM 3.0 350 70.57 54.53 58.48 59.43 30.38 32.20 50.93
RWKV-4 3.0 - 72.42 58.75 57.3 62.92 35.15 36.20 53.79
HGRN2 2.9 100 74.59 59.92 57.46 62.58 29.86 35.00 53.24
Table 7: Performances Comparison on the Long Range Arena benchmark. The results of other models come from HGRN2 outperforms HGRN1 on all datasets. ${}^{\dagger}$ indicates the results reported by Alonso et al. (2024).
Model ListOps Text Retrieval Image Pathfinder Path-X AVG.
@@ 4.3 LRA
Long Range Arena (Tay et al., 2021b) is a benchmark for accessing the model’s long-dependency modeling ability. We use HGRN1’s setting, and compare wth existing methods shown below.
Baselines.
Transformer (Tay et al., 2021b), Cosformer (Qin et al., 2022c), FLASH (Hua et al., 2022), S4 (Gu et al., 2022b), DSS (Gupta et al., 2022a), TNN (Qin et al., 2023a), S5 (Smith et al., 2023), Mega (Ma et al., 2022), SGConv (Li et al., 2022), LRU (Orvieto et al., 2023), Mamba (Gu & Dao, 2023), Griffin (De et al., 2024).
@@
Table 7 shows the result. We can see that HGRN2 outperforms HGRN1 and is competitive with other state-of-the-art models, while Mamba and Griffin failed to have a high accuracy, indicating the importance of lower bound as ablated in HGRN1.
4.4 Image modeling.
Setting
For the image classification task, we referred to the configuration of HGRN1 and train it on ImageNet-1k, also compared to TNN and the vanilla transformer
@@
Table 8 shows the result. We can see that HGRN2 outperforms HGRN1 with a similar parameter size, while having advantage over previous TNN (Qin et al., 2023a) and Deit models (Touvron et al., 2021).
Table 8: Performances comparison of image classification on ImageNet-1k. HGRN2 performs favorably than competing methods with similar parameter sizes.
@@
Model Top1 Acc Param (M) Top1 Acc Parma (M)
@@ 5 Related work
Linear recurrent models mainly include linear RNNs, state-space models, and linear attention. State-space models (SSMs) are gaining great attention since the seminal work S4 (Gu et al., 2022a) and its more efficient diagonalized version (Gu et al., 2022c). Despite excellent performance in the LRA benchmark, it has been shown to have inferior performance in language modeling. Gating mechanisms have been shown to be crucial in improving SSMs’ language modeling performance (Mehta et al., 2023; Wang et al., 2022; Gu & Dao, 2023). Gupta et al. (2022c) build the connection between SSM and linear RNN. Orvieto et al. (2023) proposes a linear RNN layer (i.e., LRU) inspired by SSMs. Peng et al. (2023) successfully scale linear RNN models to billions of parameters for the first time.
For linear attention models, their language modeling performance has been underperforming softmax attention for a long time. Several improvements have been proposed to bridge the performance gap: (i) incorporating the forgetting mechanism (Peng et al., 2021; Schlag et al., 2021; Sun et al., 2023a; Qin et al., 2023b; Yang et al., 2023), (ii) using local attention (Qin et al., 2022a; Zhang et al., 2023; Arora et al., 2024), (iii) using higher-order polynomial feature map (Arora et al., 2024; Kacham et al., 2023) to make the resulting attention distribution more sharp (Zhang et al., 2024).
@@
Gated recurrent models with matrix-valued recurrent state have been investigated in the literature of Neural Turing Machine (NTM, Graves et al. 2014) and linear Transformer (Katharopoulos et al., 2020). In NTM, the number of memory slots can be regarded as the state expansion ratio discussed in this work. NTM also included data-dependent decays in the form of erase vectors. However, NTM is hard to parallelize and thus slow to train in practice. ABC (Peng et al., 2022) could be considered a simplified and parallelizable version of NTM. The linear transformer is known to have the recurrent form (Katharopoulos et al., 2020) and is known to be closely related to fast weight programming (FWP Schlag et al. 2021). Gated FWPs have been investigated since Schlag & Schmidhuber (2017); Zhang & Zhou (2017), and have recently been revisited in Peng et al. (2021); Mao (2022); Yang et al. (2023); Katsch (2023); Pramanik et al. (2023). In particular, Yang et al. (2023) proposed a hardware-efficient training algorithm for these types of models.
@@ 6 Conclusion
In this work, we proposed HGRN2, which improves HGRN (Qin et al., 2023c) using an outer product-based state expansion mechanism inspired by linear attention, which allows for hardware-efficient training. Experiments on multiple tasks validate the advantages of HGRN2 over HGRN1. Large-scale language modeling experiments show that HGRN2 is competitive with other state-of-the-art models.

### v2

@@ 4.1 MQAR
Multi-Query Associative Recall (MQAR) (Arora et al., 2023) is an enhanced version of the synthetic induction head dataset (Fu et al., 2023), designed to test the in-context associative recall ability of subquadratic models. Arora et al. (2023) found strong correlations between MQAR accuracy and language modeling performance. Our experimental setting strictly follows the original paper [^3] https://github.com/HazyResearch/zoology. Our hyperparameter sweep included the following ranges: expansion ratio $\in\{64,128\}$ and learning rate $\in\{1e-5,5e-5,1e-4,5e-4,1e-3,5e-3,1e-2\}$.
@@
As shown in Fig. 3, HGRN2 significantly outperforms HGRN1 across various model dimensions, demonstrating the benefits of state expansion in improving memory capacity and, consequently, in-context recall ability.
Figure 3: Results on MQAR, where the x-axis represents the model dimension and the y-axis represents accuracy. The task becomes more challenging as the sequence length increases. HGRN2 outperforms HGRN1 in all scenarios.
@@ 4.2 Language modeling
4.2.1 Wikitext-103
Table 5: Results on Wikitext-103.
Model PPL (val) PPL (test) Params (M)
@@ 4.2.1 Wikitext-103
For the Wikitext-103 experiment, we followed the configuration of HGRN1 to validate the performance of 44M models against a wide range of subquadratic models: FLASH (Hua et al., 2022), 1+elu (Katharopoulos et al., 2020), Performer (Choromanski et al., 2021), cosFormer (Qin et al., 2022b), Syn(D), Syn(R) (Tay et al., 2021a), gMLP (Liu et al., 2021), S4 (Gu et al., 2022a), DSS (Gupta & Berant, 2022), RWKV-v4 (Peng et al., 2023), LRU (Orvieto et al., 2023), HGRN1 (Qin et al., 2023c), TNN (Qin et al., 2023a), and Mamba (Gu & Dao, 2023). All reported results are from our own runs under the same settings.
@@
Table 5 shows the results. HGRN2 clearly outperforms HGRN1 but slightly underperforms Mamba.
4.2.2 Slimpajama
We conducted language modeling experiments with 1.3B and 2.7B parameters on the Slimpajama dataset (Soboleva et al., 2023), using the FlashLinearAttention (Yang & Zhang, 2024) codebase for training.  [^4] Model checkpoints are available at https://huggingface.co/fla-hub. The results, shown in Table 6, demonstrate that HGRN2 consistently outperforms other competitive linear recurrent models across three model scales. This suggests that HGRN2 provides a superior parameterization compared to GLA, as both models share an identical recurrent structure.
Table 6: Slimpajama language modeling results.
Lamb. Wiki. ARCe ARCc Hella. Lamb. PIQA Wino. Avg
[…]
1.3B parameters with 100B training tokens
Transformer++ 15.3 17.1 54.1 27.1 49.3 47.0 70.3 54.9 50.5
Mamba 16.5 18.2 57.3 26.6 48.1 43.4 69.5 53.7 49.8
RetNet 15.4 17.3 57.4 27.9 50.3 44.6 71.7 51.8 50.6
GLA 15.4 17.6 55.4 27.7 49.0 46.4 69.9 54.0 50.4
HGRN2 11.8 16.9 58.1 28.1 51.8 49.4 71.4 52.3 51.9
2.7B parameters with 100B training tokens
Transformer++ 10.7 15.2 59.8 27.5 54.2 52.3 72.7 56.2 53.8
Mamba 13.6 15.9 60.7 29.8 53.9 46.4 72.8 53.9 52.9
RetNet 11.9 15.8 59.6 28.1 54.0 49.6 72.3 53.8 52.9
GLA 12.4 15.5 59.2 29.9 54.0 50.4 71.7 55.7 53.5
HGRN2 8.8 14.6 60.8 30.3 58.7 55.4 73.0 54.2 55.4
@@ 4.2.3 The Pile
We also conducted experiments on the Pile dataset. First, we trained 150M, 350M, and 1B HGRN1 and HGRN2 models for 100B tokens, and the results are shown in Table 7. We observe that HGRN2 consistently outperforms HGRN1.
Table 7: Comparison between HGRN1 and HGRN2 on Commonsense Reasoning Tasks.
Model Bn Params Bn Tokens PIQA Hella. Wino. ARC-e ARC-c OBQA AVG
@@
HGRN1 0.35 100 66.70 38.12 51.70 49.20 25.26 30.60 43.60
HGRN2 0.39 100 69.97 46.16 52.72 53.58 23.98 32.40 46.47
HGRN1 1 100 70.89 48.02 51.62 55.64 27.90 31.60 47.61
HGRN2 1 100 74.16 54.85 56.12 58.71 27.22 34.00 50.84
Next, we scaled the token horizon to 300B and trained strong baseline models, Mamba and LLaMA, under the same settings for comparison. We also compared them against several open-sourced language models, such as OPT (Zhang et al., 2022), Pythia (Biderman et al., 2023), BLOOM (Scao et al., 2022), and RWKV-4 (Peng et al., 2023). We found that HGRN2 performs competitively with Mamba, LLaMA, and other open-sourced LLMs.
Table 8: Comparison between HGRN2 and other open-sourced language models, alongside strong baseline models (LLaMA and Mamba re-trained under the same settings), on Commonsense Reasoning Tasks. indicates our own trained model.
Model Bn Params Bn Token PIQA Hella. Wino. ARC-e ARC-c OBQA AVG
@@
BLOOM 0.56 350 64.09 46.97 52.80 47.35 29.38 28.20 42.23
RWKV-4 0.43 - 67.52 39.00 51.14 52.86 25.17 32.40 45.00
Llama 0.4 350 67.19 38.75 52.19 49.24 23.72 30.00 43.51
Mamba 0.4 300 67.90 40.74 52.72 53.07 24.74 31.20 45.06
HGRN2 0.4 300 67.74 40.32 51.78 54.21 24.83 31.20 45.01
@@
Llama 1.0 300 69.97 47.04 52.72 57.07 26.18 32.60 47.93
Mamba 1.0 300 71.27 50.15 56.35 58.71 29.27 31.20 49.45
HGRN2 1.0 300 71.65 49.52 54.38 60.27 28.07 33.40 49.55
OPT 2.7 300 73.83 60.60 61.01 60.77 31.31 35.20 53.79
@@
BLOOM 3.0 350 70.57 54.53 58.49 59.43 30.38 32.20 50.77
RWKV-4 3.0 - 72.42 58.75 57.30 62.92 35.15 36.20 53.79
Llama 3.0 350 73.18 57.88 59.59 63.93 33.51 35.40 53.93
Mamba 3.0 300 74.92 61.68 59.19 65.33 31.45 35.60 55.31
HGRN2 3.0 300 74.10 61.48 58.64 65.61 34.47 35.60 54.98
Llama 7.0 300 75.19 64.39 61.88 67.55 35.41 35.00 56.57
HGRN2 7.0 300 76.50 66.96 61.40 69.02 36.86 38.00 58.12
To evaluate long-context abilities, we conducted tests on SCROLLs (Shaham et al., 2022) and found that HGRN2 exhibits better scaling behavior compared to Mamba, indicating stronger long-context capabilities, potentially due to its larger recurrent state size. However, we also observed that the 7B HGRN2 model is still not as strong as the LLaMA model, suggesting that the scaling behavior of linear models for long-context modeling remains an area for further study.
Table 9: Performance Comparison on SCROLLS. R-1/2/L stand for parameter size, tokens, and rouge-1/rouge-2/rouge-l, respectively.
Model Params Token GovRep SumScr QMSum Qspr Nrtv QALT CNLI Avg $\uparrow$
Bn Bn R-1/2/L R-1/2/L R-1/2/L F1 F1 EM EM
Llama 0.4 300 8.2/3.5/6.2 11.3/1.6/8.7 10.7/2.1/9.4 17.8 15.4 28.0 13.9 10.5
Mamba 0.4 300 8.2/2.4/6.2 11.2/1.8/8.9 9.3/1.6/8.4 14.9 11.6 25.8 19.4 10.0
HGRN2 0.4 300 15.3/3.5/10.9 7.4/0.8/6.2 8.3/1.2/7.4 12.4 10.9 26.4 31.5 10.9
Llama 1.0 300 12.9/3.1/9.4 9.5/0.8/7.7 10.9/2.2/9.4 22.8 16.0 28.4 9.9 11.0
Mamba 1.0 300 15.2/4.2/10.6 12.3/1.6/9.4 13.9/3.1/11.7 18.3 14.7 26.7 9.1 11.6
HGRN2 1.0 300 14.9/4.2/10.5 11.4/1.4/9.2 10.9/2.3/9.7 16.2 15.1 27.8 10.6 11.1
Llama 3.0 300 11.2/4.9/8.1 11.9/1.9/9.3 16.1/4.3/12.9 28.6 20.8 30.4 20.2 13.9
Mamba 3.0 300 21.5/6.6/13.9 13.2/2.0/10.1 15.0/3.2/12.3 22.1 17.9 28.8 24.0 14.7
HGRN2 3.0 300 21.7/6.6/14.1 14.6/2.1/10.8 12.5/2.7/10.6 25.4 18.8 28.9 31.9 15.4
Llama 7.0 300 17.4/7.3/11.4 12.9/1.8/10.0 14.6/3.7/11.8 32.4 22.3 33.8 10.0 14.6
HGRN2 7.0 300 14.9/5.2/10.2 15.4/2.4/11.1 14.3/3.0/11.8 27.1 19.6 30.1 10.0 13.5
To test the retrieval ability of our trained 3B models, we ran the easy mode of the Needle in a Haystack Test.  [^5] In this mode (Shen, 2024; Shen et al., 2024), both the question and answer (QA pair) are embedded within a lengthy text, challenging the model to locate and respond to the query. This mode is particularly suitable for base models without instruction tuning. In contrast, the standard mode only places the answer within the long context, requiring the model to understand the question and find the relevant answer. LLaMA almost achieves perfect retrieval performance for evaluation lengths no greater than the training length. As shown in Figure 4, HGRN2 and Mamba still face difficulties in retrieval tasks; however, HGRN2 outperforms Mamba due to its larger state size, enabled by linear attention-styled state expansion.
Figure 4: Easy mode Needle in a Haystack Test on 3B models: Mamba (left) and HGRN2 (right). The evaluation context length is 16K, and the models were trained on a sequence length of 8K.
4.3 Long Range Arena
Table 10: Results on LRA. indicates the results reported by Alonso et al. (2024).
Model ListOps Text Retrieval Image Pathfinder Path-X AVG
@@ 4.3 Long Range Arena
Long Range Arena (Tay et al., 2021b) is a benchmark designed to assess a model’s ability to handle long-range dependencies. We used HGRN1’s configuration and compared it with existing methods, as shown below.
@@
Table 10 shows the results. HGRN2 outperforms HGRN1, while Mamba and Griffin failed to achieve high accuracy on this benchmark.
4.4 Image Modeling
Setting.
For the image classification task, we followed the configuration of HGRN1 and trained it on ImageNet-1k, comparing it with TNN and the vanilla transformer.
@@ 4.4 Image Modeling
Table 11 shows the results. HGRN2 outperforms HGRN1 with a similar parameter size, while also demonstrating an advantage over previous TNN (Qin et al., 2023a) and DeiT models (Touvron et al., 2021).
Table 11: Performances comparison of image classification on ImageNet-1k. HGRN2 performs favorably compared to competing methods with similar parameter sizes.
@@
Model Top-1 Acc Params (M) Top-1 Acc Params (M)
@@ 5 Related work
Linear recurrent models mainly include linear RNNs, state-space models, and linear attention. State-space models (SSMs) are gaining great attention since the seminal work S4 (Gu et al., 2022a) and its more efficient diagonalized version (Gu et al., 2022b). Despite excellent performance in the LRA benchmark, it has been shown to have inferior performance in language modeling. Gating mechanisms have been shown to be crucial in improving SSMs’ language modeling performance (Mehta et al., 2023; Wang et al., 2022; Gu & Dao, 2023). Gupta et al. (2022) build the connection between SSM and linear RNN. Orvieto et al. (2023) proposes a linear RNN layer (i.e., LRU) inspired by SSMs. Peng et al. (2023) successfully scale linear RNN models to billions of parameters for the first time.
For linear attention models, their language modeling performance has been underperforming softmax attention for a long time. Several improvements have been proposed to bridge the performance gap: (i) incorporating the forgetting mechanism (Peng et al., 2021; Schlag et al., 2021; Sun et al., 2023; Qin et al., 2023b; Yang et al., 2023; Peng et al., 2024), (ii) using local attention (Qin et al., 2022a; Zhang et al., 2023; Arora et al., 2024; Ren et al., 2024), (iii) using higher-order polynomial feature map (Arora et al., 2024; Kacham et al., 2023) to make the resulting attention distribution more sharp (Zhang et al., 2024), (iv) using more expressive yet efficient recurrent update rule (Schlag et al., 2021; Yang et al., 2024; Liu et al., 2024; Sun et al., 2024a).
@@
Gated recurrent models with matrix-valued recurrent state have been investigated in the literature of Neural Turing Machine (NTM Graves et al. 2014) and linear Transformer (Katharopoulos et al., 2020). In NTM, the number of memory slots can be regarded as the state expansion ratio discussed in this work. NTM also included data-dependent decays in the form of erase vectors. However, NTM is hard to parallelize and thus slow to train in practice. The linear transformer is known to have the recurrent form (Katharopoulos et al., 2020) and is known to be closely related to fast weight programming (FWP Schlag et al. 2021). Gated FWPs have been investigated since Schlag & Schmidhuber (2017); Zhang & Zhou (2017), and have recently been revisited in Peng et al. (2021); Mao (2022); Yang et al. (2023); Katsch (2023); Pramanik et al. (2023). In particular, Yang et al. (2023) proposed a hardware-efficient training algorithm for these types of models.
More recently, Mamba2 (Dao & Gu, 2024), xLSTM (Beck et al., 2024), and Gated Retention (Sun et al., 2024b) have shown that sharing data-dependent decays across different dimensions within the same head is effective. This approach improves efficiency over GLA because intra-chunk computations are more amenable to tensor core-based matrix multiplication acceleration, at the cost of sacrificing the fine-grainedness of decays. In GLA/HGRN2, each head dimension has its own decay rate, whereas in Mamba2/xLSTM/Gated Retention, all dimensions share the decay under a single head. It is an interesting question to study how much improvement fine-grained decay will bring.
@@ 6 Conclusion
In this work, we propose HGRN2, an enhancement of HGRN (Qin et al., 2023c) using an outer product-based state expansion mechanism inspired by linear attention, enabling efficient training. Experiments across multiple tasks validate the advantages of HGRN2 over HGRN1.
Acknowledgement
We thank Yu Zhang for conducting some language modeling experiments and for the valuable discussions.